The head of the mathematics department announced that the mean score of Grade 11 students in the second periodic test in statistics was 89, And a standard deviation was 12. One student believed that the main score was less than this, So the student randomly selected 34 students and computer do you mean score, And obtained I mean score of 85. At 0.01 level of significance, Construct the critical regions.


1
Expert's answer
2022-05-30T23:15:40-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=89"

"H_a:\mu<89"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a left-tailed test is (using t-table) "z_c = -2.3263."

The rejection region for this left-tailed test is "R = \{z: z < -2.3263\}."

The z-statistic is computed as follows:





"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{85-89}{12/\sqrt{34}}\approx-1.94"

Since it is observed that "z = -1.94> -2.3263=z_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 89, at the "\alpha = 0.01" significance level.